OpenStudy (anonymous):

4e^(X^2)=800 solve for x satellite where are you?

OpenStudy (anonymous):

right here

OpenStudy (anonymous):

divide by 4

OpenStudy (anonymous):

\[e^{x^2}=200\]

OpenStudy (anonymous):

rewrite in equivalent logarithmic form

OpenStudy (anonymous):

\[x^2=\ln(200)\]

OpenStudy (anonymous):

solve for x: \[x=\pm\sqrt{\ln200}\]

OpenStudy (anonymous):

also note that \[ln(200)=ln(2\times 100)=ln(2)+ln(100)=ln(2)+ln(10^2)\] \[=ln(2)+2ln(10)\]

OpenStudy (anonymous):

not that this helps much

OpenStudy (anonymous):

k?

OpenStudy (anonymous):

for sure? i had x^2lne=ln200.....then x^2+1=ln 200.....is that wrong?

OpenStudy (anonymous):

ok but \[x^2\ln(e)=x^2\times 1 =x^2\]

OpenStudy (anonymous):

not \[x^2+1\]

OpenStudy (anonymous):

ohh......okay....i see why that was wrong. you are amazing ;)

OpenStudy (anonymous):

don't take that step in any case. remember that \[e^x=y\] says exactly the same thing as \[y=\ln(x)\]

OpenStudy (anonymous):

you are taking too many steps. write in "equivalent logarithmic form" and you will not have to mess with \[\ln(e)\]

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

welcome!

OpenStudy (anonymous):

i gotta run, do you have another quick one?

OpenStudy (anonymous):

convert log$3(22) into natural. is it ln22/ln3?

OpenStudy (anonymous):

yes change of base formula you have it.

OpenStudy (anonymous):

okay bye.wish i could put you in my pocket. hope to find you again